OpenStudy (anonymous):

Allen bought 20 stamps at the post office in 37¢ and 20¢ denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps did Allen buy? Set up a system of equations. Then solve using substitution or elimination.

hero (hero):

x + y = 20 .37x + .20y = 7.06

hero (hero):

Those are your two equations

OpenStudy (anonymous):

let a= number of stamps costing .37 and b=number of stamps costing .20 a+b=20 .37a+.20b=7.06

OpenStudy (anonymous):

use substitution a+b=20 a=20-b

OpenStudy (anonymous):

.37(20-b)+.20b=7.06 7.4-.37b+.20b=7.06 7.4-.17b=7.06 -.17b=-.34 b=2

OpenStudy (anonymous):

plug b back into one of the equations and solve for a. 2+a=20 a=18 he bought 18 stamps that cost 37 cents and 2 stamps that cost 20 cents

OpenStudy (gw2011):

Let x=37 cent stamps Let y=20 cent stamps x+y=20 0.37x+0.20y=7.06 By using substitution you get: y=20-x 0.37x+0.20(20-x)=7.06 0.37x+4-0.20x=7.06 0.17x=3.06 x=18 y=20-18 y=2 Therefore, there are 18 stamps at 37 cents and 2 stamps at 20 cents